Mardi recieved an inheritance of $60000. She invested part at 7% and deposited the remainder in tax-free bonds at 11%. Her total annual income from the investments was $5400. Find the amount invested at 7%.
Answer by mananth(16949)

Total invested 60,000
Interest----- 5400

Part I 7.00% per annum ---x
Tax free bonds 11.00% per annum ---y
Total investment
x + y= 60000 -------------1
Interest on both investments
7.00% x + 11.00% y= 5400
Multiply by 2
7 x + 11 y= 540000.00 --------2
Multiply (1) by -7
we get
-7 x -7 y= -420000.00
Add this to (2)
0 x 4 y= 120000
divide by 4
y = 30000
Part I 7.00% $ 30000
Tax free bonds 11.00% $ 30000


CHECK
30000 --------- 7.00% ------- 2100.00
30000 ------------- 11.00% ------- 3300.00
Total -------------------- 5400.00

m.ananth@hotmail.ca